### v2

@@ 4.2.5 Distance between messages
The final metric to be considered in our analysis is the distance between two successive AIS messages. The speed filtering process described in Section 3.2 involves removing trajectory segments where the SOG falls below one knot. This exclusion results in significant spatial gaps between successive messages within a trajectory and potentially leads to substantial temporal gaps if the vessel continues sailing. If the vessel increases its SOG to over one knot, its messages appear again in the data. In these cases, the discrepancy between the calculated speed (derived from positional and temporal data) and the average reported speed might be insufficiently large to be identified as split-points, based on the $(1-\alpha)\cdot 100=95$ percentile criterion.
To additionally capture these special cases, we also label two consecutive messages a split-point if their distance in miles, as calculated via the Haversine formula, is greater than the $(1-\alpha)\cdot 100=95$ percentile of the distance distribution obtained for the month of June 2021, which sets the boundary to $1.17nm$, with $nm$ standing for nautical miles.
@@ 4.3 Re-joining trajectories
Data: Trajectory to split $\mathcal{T}^{>}$, Empirical quantile functions for change in speed $\hat{q}^{\Delta_{SOG}}$, tuning rate $\hat{q}^{{ROT}}$, time difference $\hat{q}^{\Delta_{t}}$, distance $\hat{q}^{dist}$, and reported against calculated speed $\hat{q}^{\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu-\widehat{SOG}}$, quantile threshold $\alpha$ (0.05 for this study).
@@
$s\leftarrow\hat{q}^{\Delta_{SOG}}(1-\alpha)$
$r\leftarrow[\hat{q}^{{ROT}}(\alpha/2),\hat{q}^{{ROT}}(1-\alpha/2)]$
$t\leftarrow\hat{q}^{\Delta_{t}}(1-\alpha)$
$d\leftarrow\hat{q}^{dist}(1-\alpha)$
[…]
@@
                 or
             $\frac{\operatorname{cog}(m_{i+1})-\operatorname{cog}(m_{i})}{\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})}\notin r$ (Section 4.2.3)
                 or
             $\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})>t$ (Section 4.2.1)
                 or
             $\operatorname{hav}(m_{i+1},m_{i})>d$ (Section 4.2.5)
                 or
             $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}-\widehat{SOG}_{m_{i}}^{m_{i+1}}\notin b$ (Section 4.2.4)
@@
4.4 Applications of the split-point procedure to data
To demonstrate the effectiveness of the proposed split-point methodology, Figure 11 showcases a collection of trajectories derived from segments of the Little Belt and harbors of Aabenraa and Assens in Denmark, both with and without the application of the split-point procedure. In the absence of this procedure (as depicted in Figure 11(a)), numerous trajectories exhibit anomalously large distances between consecutive messages. This results in unrealistic jumps in the trajectory paths, leading to instances where the trajectories erroneously appear to traverse over land. This behavior can arise if multiple vessels transmit messages from a single MMSI, which is a common occurrence (Wu et al., 2017; Zhao et al., 2018; Yan et al., 2020), if trajectories are interrupted by the SOG-filter constraint or if vessels deliberately disable their AIS transceiver and re-enable it at a different location. In the filtered version, as depicted in Figure 11(b), the application of the split-point procedure effectively eliminates these anomalies, thereby accurately reconstructing the underlying structure of the waterway network. Quantitative information about the procedure can be found in Figure 11(c).
The split-point procedure also allows for significant insights into the navigational behaviors of vessels across different regions and types. This technique is particularly effective in generating ship-type dependent density maps, which illustrate the varying concentrations of vessel traffic along established routes and can also be used to discover new routes. This procedure allows key navigational routes and traffic flow dynamics specific to various ship types to be analyzed effectively. The methodology’s utility and effectiveness are demonstrated in Figures 14 and 15, which present detailed density maps of vessel traffic routes in the area of this research. Detailed information for practitioners on how the accompanying Python package implements Algorithm 1 and is utilized to generate these results is described in C.
@@ 4.4 Applications of the split-point procedure to data
(b) Trajectories after applying the split-point procedure from Section 4.2.
Number of Raw Messages 10 420 4091042040910\,420\,40910 420 409 Percentage of Duplicates 3.11% Number of Trajectories 672 803672803672\,803672 803
Number of Split-Points 2 257 02322570232\,257\,0232 257 023 Number of Rejoined Tracks 47 4924749247\,49247 492 Average Trajectory length $2.66$ $nm$
@@
Figure 11: Effects of the split-point filter applied to an area around the Little Belt in Denmark. Only the time frame from 10/07/21 - 19/07/21 has been used to avoid clutter.
5 Time scale of empirical distributions
The split-point methodology outlined in Section 4.2 requires the computation of five distinct empirical distributions, using their quantiles to establish thresholds. In this process, we explicitly used trajectory data from one month of AIS records as the basis for each distribution. Figure 12 displays the quantile functions for all metrics, each calculated using different amounts of trajectory data (given as days and months). Visually, these distributions appear somewhat similar, leading to the assumption that the data follows identical distributions over various time horizons.
The computed quantile values in Table 4 only show minor differences in absolute terms, which leaves leeway to the choice of the amount of data used to calculate the quantiles. However, due to the huge sample sizes (over $10^{6}$ observations) that lead to very narrow confidence intervals, almost all quantiles are tested statistically different. Ultimately, we use one month of data throughout the study as a practical trade-off between accuracy and computational burden.
6 Spatial properties of split trajectories
@@
We have formulated various trajectories, each comprising a reliable and consistent series of messages, notably free from anomalies. Nonetheless, we have no tool to assess the spatial properties of the extracted trajectories, especially concerning our original aim of extracting long, clean, and uninterrupted trajectories.
@@
$\bar{c}(\mathcal{T}_{k}^{>})=|\mathcal{T}_{k}^{>}|^{-1}\sum_{m\in\mathcal{T}_{k}^{>}}\frac{\mathbf{p}^{\top}\mathbf{q}}{\|\mathbf{p}\|_{2}\|\mathbf{q}\|_{2}}$ (6)
@@
$\mathbf{p}=\begin{bmatrix}\operatorname{lat}(m_{i})-\operatorname{lat}(m_{i-1})\\ \operatorname{lon}(m_{i})-\operatorname{lon}(m_{i-1})\end{bmatrix},\quad\mathbf{q}=\begin{bmatrix}\operatorname{lat}(m_{i+1})-\operatorname{lat}(m_{i})\\ \operatorname{lon}(m_{i+1})-\operatorname{lon}(m_{i})\end{bmatrix}$
@@
$\bar{\Delta}(\mathcal{T}^{>}_{k})=\pi^{-1}\cos^{-1}[\bar{c}(\mathcal{T}^{>}_{k})]\times 180\quad\text{for}|\mathcal{T}^{>}_{k}|>3$ (7)
@@ 7 Discussion
This study provides a framework for trajectory extraction from raw, encoded AIVDM sentences. A multi-level filtering and splitting procedure has been proposed to construct long, clean, uninterrupted trajectories. Particular emphasis lies on the data-driven nature of the splitting process of this study. While other approaches (Sang et al., 2015; Zhao et al., 2018; Zhang et al., 2018; Yuan et al., 2019; Chen et al., 2020), too, propose functioning extraction processes, many threshold values used in those studies are still expert-derived, questioning their generalizability beyond the locally tested data. This study, instead, proposes to generate an empirical cumulative distribution function for the metric under consideration and use its 95th percentile bounds to determine the threshold values.
The methodology developed in this study demonstrates effective performance and aligns to establish a self-explanatory, data-driven approach for trajectory extraction. However, it is acknowledged that the decision to utilize the 95th percentile of the distributions as threshold bounds may be perceived as somewhat arbitrary. This choice, while not without its rationale, invites further research, as many derived thresholds have no specific meaning regarding established AIS transceiving regulations or the maritime domain in general.
@@ 8 Conclusion
Maritime AIS data is inherently noisy due to technical inaccuracies, seafarer inattention, or lacking compliance. However, trajectories extracted from AIS messages are valuable for maritime surveillance, domain awareness, and algorithmic testing, for which we propose a data-driven, open-source trajectory extraction and refinement framework. The extraction process is based on a split-point procedure that uses quantile boundaries of five empirically derived distribution functions to decide whether two consecutive messages belong to the same trajectory or should be split. We found that this procedure effectively eliminates anomalies in the data without resorting to expert-derived thresholds, making it universally adaptable to any AIS data set. Nevertheless, carefully scrutinizing the quantile boundary and threshold values is recommended, as they may vary significantly depending on data availability. For spatial understanding, we introduced the concept of average absolute change of course, $\bar{\Delta}(\cdot)$, which allows us to assess the presence or absence of erratic positional movement in a trajectory and recommended the number of messages per trajectory ($n_{\mathrm{msg}}$) as well as the convex hull area ($A^{C}$) as means to control $\bar{\Delta}(\cdot)$. The entire extraction and assessment pipeline, described by this article, is publicly available as an open-source Python package at Paulig (2024).

### v3

@@ 6.2.5 Distance between messages
The final metric to be considered in our analysis is the distance between two successive AIS messages. The speed filtering process described in Section 5.2 involves removing trajectory segments where the SOG falls below one knot. This exclusion results in significant spatial gaps between successive messages within a trajectory, potentially leading to substantial temporal gaps if the vessel continues sailing. If the vessel increases its SOG to over one knot, its messages appear again in the data. In these cases, the discrepancy between the calculated speed (derived from positional and temporal data) and the average reported speed might be insufficiently large to be identified as split points based on their percentile criterion.
To additionally capture these special cases, we also label two consecutive messages a split-point if their distance in miles, as calculated via the Haversine formula, is greater than the $(1-\alpha)$ quantile of the distance distribution obtained for the year 2021, which are depicted in Figure 6.
Length bin $\alpha$ Length bin $\alpha$ Length bin $\alpha$
0.1 0.05 0.01 0.1 0.05 0.01 0.1 0.05 0.01
$[0,25)m$ 0.727 1.001 1.969 $[75,100)m$ 0.922 1.095 1.352 $[150,175)m$ 1.148 1.424 1.879
$[25,50)m$ 0.771 1.004 1.574 $[100,125)m$ 0.992 1.205 1.517 $[175,200)m$ 1.162 1.391 1.775
$[50,75)m$ 0.895 1.072 1.332 $[125,150)m$ 1.160 1.409 1.808 $[200,\infty)m$ 1.177 1.444 1.994
Table 6: Cutoff quantile values for the distance of two consecutive messages in miles for selected values of $\alpha$ across the different length intervals.
@@ 6.3 Re-joining trajectories
Data: Trajectory to split $\mathcal{T}^{>}$, Ship length $l$, Length-dependent empirical quantile functions for change in speed $\hat{q}_{l}^{\Delta_{SOG}}$, tuning rate $\hat{q}_{l}^{{ROT}}$, time difference $\hat{q}_{l}^{\Delta_{t}}$, distance $\hat{q}_{l}^{dist}$, and reported against calculated speed $\hat{q}_{l}^{\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu-\widehat{SOG}}$, quantile threshold $\alpha$.
@@
$s\leftarrow\hat{q}_{l}^{\Delta_{SOG}}(1-\alpha)$
$r\leftarrow[\hat{q}_{l}^{{ROT}}(\alpha/2),\hat{q}_{l}^{{ROT}}(1-\alpha/2)]$
$t\leftarrow\hat{q}_{l}^{\Delta_{t}}(1-\alpha)$
$d\leftarrow\hat{q}_{l}^{dist}(1-\alpha)$
$b\leftarrow[\hat{q}_{l}^{\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu-\widehat{SOG}}(\alpha/2),\hat{q}_{l}^{\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu-\widehat{SOG}}(1-\alpha/2)]$
@@
                  or $\frac{\operatorname{cog}(m_{i+1})-\operatorname{cog}(m_{i})}{\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})}\notin r$ (Section 6.2.3)
                  or $\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})>t$ (Section 6.2.1)
                  or $\operatorname{hav}(m_{i+1},m_{i})>d$ (Section 6.2.5)
                  or $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}-\widehat{SOG}_{m_{i}}^{m_{i+1}}\notin b$ (Section 6.2.4)
@@
7 Applications of the split-point procedure to data
To demonstrate the effectiveness of the proposed split-point methodology, Figure 12 showcases a collection of trajectories derived from segments of the Øresund and harbors of Copenhagen in Denmark and Malmö and Helsingborg in Sweden, both with and without the application of the split-point procedure. In the absence of this procedure (as depicted in Figure 12(a)), numerous trajectories exhibit anomalously large distances between consecutive messages. This results in unrealistic jumps in the trajectory paths, leading to instances where the trajectories erroneously appear to traverse over land. This behavior can arise if multiple vessels transmit messages from a single MMSI, which is a common occurrence (Wu et al., 2017; Zhao et al., 2018; Yan et al., 2020), if trajectories are interrupted by the SOG-filter constraint or if vessels deliberately disable their AIS transceiver and re-enable it at a different location. In the filtered version, as depicted in Figure 12(b), the application of the split-point procedure effectively eliminates these anomalies, thereby accurately reconstructing the underlying structure of the waterway network and, due to the different coloring by the length of the ship, one can also gain valuable insights into the distribution of ship sizes along the waterway, enabling a more comprehensive analysis of traffic patterns and the potential identification of areas with specific navigational challenges or congestion points based on vessel dimensions.Quantitative information about the procedure can be found in Figure 12(c).
The split-point procedure also allows for significant insights into the navigational behaviors of vessels across different regions and types. This technique is particularly effective in generating ship-type dependent density maps, which illustrate the varying concentrations of vessel traffic along established routes and can also be used to discover new routes. This procedure allows key navigational routes and traffic flow dynamics specific to various ship types to be analyzed effectively. The methodology’s utility and effectiveness are demonstrated in the Appendix B in Figures 16 and 17, which present detailed density maps of vessel traffic routes in the area of this research.
On an individual trajectory level Figure 13 demonstrates the effects of choosing different values of $\alpha$ on an unfiltered trajectory by showing how different values of $\alpha$ impact the number of splits in the trajectory. As $\alpha$ decreases, the trajectory becomes less interrupted with fewer splits. This demonstrates the trade-off between the level of data retention and the degree of realism, highlighting how lower $\alpha$ values result in a more generalized path with increasingly unreliable jumps in the metrics between two consecutive messages. The properties of trajectories extracted with a given value of $\alpha$ are easily interpretable by guaranteeing that all consecutive messages within the trajectory lie inside the navigational bounds of $(1-\alpha)\cdot 100\%$ of all vessels used to determine the thresholds. This ensures that the extracted trajectories represent common navigational patterns, making them reliable and consistent with the overall traffic flow while filtering out outliers or unusual movements.
@@ 7 Applications of the split-point procedure to data
(b) Trajectories after applying the split-point procedure from Section 6.2 with $\alpha=0.03$.
Number of Raw Messages 26 612 0472661204726\,612\,04726 612 047 Percentage of Duplicates 0.46% Number of Trajectories 722 624722624722\,624722 624
Number of Split-Points 1 444 72414447241\,444\,7241 444 724 Number of Rejoined Tracks 20 5582055820\,55820 558 Average Trajectory length $4.53$ $nm$
@@
Figure 12: Effects of the split-point filter applied to an area in the Øresund around Copenhagen in Denmark. Data from 10/08/2024 - 19/08/24 is used.
Moreover, due to its data-driven nature, the split-point methodology is highly generalizable and can be effectively applied to any geographical region. Whether dealing with complex networks of inland waterways, coastal regions, or open ocean environments, the underlying principles of the split-point procedure remain consistent and adaptable. The flexibility of this approach allows for its application in diverse maritime contexts, making it a valuable tool for global maritime navigation analysis but also as a precursor for algorithmic testing procedures that rely on realistic real-world trajectories. Practitioners can leverage the accompanying Python package, as detailed in Appendix C, to use and adapt Algorithm 1 and generate results similar to those presented, regardless of the geographical region under study. This generalizability ensures that the insights gained from this methodology are not confined to the specific regions of the North Sea but are applicable worldwide.
Figure 13: Results of our proposed split-point procedure under different values of $\alpha$, where lower values split less aggressively at the cost of larger margins of accepted values between messages. Differently colored parts of a trajectory indicate sub-trajectories. On the rightmost illustration, the lower-right part of the trajectory gets filtered out entirely as it was decomposed into several single-message trajectories, which automatically get discarded. The most significant positional gap in the original trajectory is $3.11nm$.
8 Comparison
When comparing trajectory extraction processes from AIS data, it is essential to recognize the inherent limitations in quantitatively evaluating the performance of different methods. As mentioned in the introduction, AIS data, while a valuable source of real-time maritime information, is subject to various errors and inconsistencies. These can include inaccuracies in positional data due to GPS errors, data loss from signal gaps, or intentional manipulation of AIS transmissions. Consequently, no definitive ground truth can be used as a baseline for comparison.
Given this lack of an absolute reference, direct quantitative comparisons of trajectory extraction methods become problematic. Traditional metrics like accuracy or error rates, which rely on a known correct answer, are not applicable in this context. Instead, comparisons between methods must often be based on qualitative assessments or indirect metrics that reflect the performance of the extraction processes under realistic conditions.
Considering the above restrictions, we compare our approach to two commonly used extraction methods, namely the ones by Zhao et al. (2018) and Guo et al. (2021). Both are popular approaches, whereby Guo et al. (2021) is a pure exclusion process that does not split trajectories but removes erroneous messages. Zhao et al. (2018), on the other hand, acts as the base of our proposed framework, as it also splits trajectories. It is to be mentioned that the approach from Guo et al. (2021) also features an extended trajectory interpolation scheme, which we will not implement in our comparison because neither the approach of Zhao et al. (2018) nor ours interpolate filtered trajectories, rendering the comparison skewed.
For our comparison, we will investigate the maximum jump in metrics such as time, distance, and velocity inside each trajectory, examine the values, and view them in a maritime context. For all compared methods we start with a time-sorted collection of AIS messages coming from one MMSI.
Setup
Zhao et al. (2018) uses three steps to filter messages. (1) All messages whose absolute difference between receiving time and sending time is more than $5s$. (2) For a pair of consecutive messages $m_{1},m_{2}$, inside a trajectory, a split point is defined if $\widehat{SOG}_{m_{1}}^{m_{2}}>15kn$ or $\operatorname{unix}(m_{2})-\operatorname{unix}(m_{1})>600s$. (3) All sub-tracks constructed in (2) are rejoined similarly as described in Section 6.3, with the exception that only the condition $\widehat{SOG}_{m_{1}}^{m_{2}}>15kn$ is taken for a decision.
The exclusion process of Guo et al. (2021) consists of two steps. (1) Turning rates are calculated in the same way as in Section 6.2.3, and a turning rate limit $\Delta c_{lim}$ is defined. If the turning rate of any two messages, $m_{1}$ and $m_{2}$, is higher than the limit, the message $m_{2}$ is discarded. (2) If for any two messages, $\widehat{SOG}_{m_{1}}^{m_{2}}>v_{lim}kn$, for a velocity limit $v_{lim}$, message $m_{2}$ is discarded. Unfortunately, the authors did not provide specific values for $\Delta c_{lim}$ and $v_{lim}$ but rather reference the ship’s maneuverability under consideration and the speed limit in the area.
Example trajectory
A 223-messages-long trajectory with distinct jumps driven by a $19m$ long fishing vessel, roughly $80km$, off the coast of the Netherlands, is investigated. The grey line in Figure 14 shows the raw trajectory. Since there is no speed limit this far off the coast, we set $v_{lim}=30kn$, and for $c_{lim}=2^{\circ}$ which is a bit more the 99-percentile of the turning rate distribution for vessels with lengths $[0,25)m$ as calculated in Section 6.2.3.
Figure 14: Method comparison for different trajectory extraction approaches.
Visually comparing the different approaches in Figure 14 reveals that no other approach was able to either effectively filter out or separate the trajectory at the two large spatial jumps that span distances of $37.85nm$ and $24.15nm$ each. As seen in Table 7, the approach of Guo et al. (2021) discarded a single message or $0.45\%$ of the messages, the one of Zhao et al. (2018) discarded $54.71\%$ of the messages mainly due to mismatching sending and receiving timestamps. No splitting was performed. Guo et al. (2021) additionally allowed for a quite extensive turning rate of $5.10^{\circ}/s$.
# of discarded
messages
# of split
points
max. turning
rate $[{}^{\circ}/s]$
max. velocity
change $[kn/s]$
max. distance
$[nm]$
Zhao et al. (2018) 1 0 0.58 0.04 53.06
Guo et al. (2021) 122 - 5.10 0.10 37.87
Proposed 2 4 0.4 0.04 1.57
Table 7: Maximum values attained for consecutive messages inside a trajectory or any of its split sub-trajectories.
In contrast, the proposed method demonstrates a balanced approach with minimal message loss and more optimal spatial filtering. As highlighted in Table 7, the proposed approach discarded just two messages, introduced four split points to avoid large spatial jumps, and maintained a realistic maximum turning rate of $0.4^{\circ}/s$. These results indicate that the proposed method achieves accurate and efficient trajectory filtering, ensuring superior spatial consistency and precision.
While this approach demonstrates superior performance in the specific scenario analyzed, it is important to acknowledge that there may be better choices for some applications. The setup process for the proposed method is considerably more complex and time-consuming compared to the more straightforward approaches of Zhao et al. (2018) and Guo et al. (2021). This complexity arises from the need to pre-compute all threshold values from large amounts of data, which may require substantial computational resources. For scenarios where quick deployment or lower computational overhead is prioritized, the simplicity and ease of setup offered by other methods may outweigh the benefits of our more intricate approach. Thus, while the proposed method achieves high accuracy and precision, its applicability may be limited in contexts with critical ease of use and rapid implementation. Consequently, both methods are also implemented in the accompanying software package.
9 Spatial properties of split trajectories ( )
@@
From the approach in Section 6.2 we extracted trajectories, each comprising a reliable and consistent series of messages, notably free from anomalies. Nonetheless, we have no tool to assess the spatial properties of the extracted trajectories, especially concerning our original aim of extracting long, clean, and uninterrupted trajectories.
@@
$\bar{c}(\mathcal{T}_{k}^{>})=|\mathcal{T}_{k}^{>}|^{-1}\sum_{m\in\mathcal{T}_{k}^{>}}\frac{\mathbf{p}^{\top}\mathbf{q}}{\|\mathbf{p}\|_{2}\|\mathbf{q}\|_{2}},$ (6)
@@
$\mathbf{p}=\begin{bmatrix}\operatorname{lat}(m_{i})-\operatorname{lat}(m_{i-1})\\ \operatorname{lon}(m_{i})-\operatorname{lon}(m_{i-1})\end{bmatrix},\quad\mathbf{q}=\begin{bmatrix}\operatorname{lat}(m_{i+1})-\operatorname{lat}(m_{i})\\ \operatorname{lon}(m_{i+1})-\operatorname{lon}(m_{i})\end{bmatrix},$
@@
$\bar{\Delta}(\mathcal{T}^{>}_{k})=\pi^{-1}\cos^{-1}[\bar{c}(\mathcal{T}^{>}_{k})]\times 180,\quad\text{for}\quad|\mathcal{T}^{>}_{k}|>3,$ (7)
@@ 10 Discussion
This study provides a framework for trajectory extraction from raw, encoded AIVDM sentences. A multi-level filtering and splitting procedure has been proposed to construct long, clean, uninterrupted trajectories. Particular emphasis lies on the data-driven nature of the splitting process of this study. While other approaches (Sang et al., 2015; Zhao et al., 2018; Zhang et al., 2018; Yuan et al., 2019; Chen et al., 2020), too, propose functioning extraction processes, many threshold values used in those studies are still expert-derived, questioning their generalizability beyond the locally tested data. This study, instead, proposes to generate an empirical cumulative distribution function for the metric under consideration and control the statistical properties of the extracted trajectories freely by using the $1-\alpha$ quantile bounds to determine the threshold values.
The methodology developed in this study demonstrates effective performance and aligns to establish a self-explanatory, data-driven approach for trajectory extraction. However, it is acknowledged that there exists no specific guidance on the selection of the quantile parameter $\alpha$, thus the final decision must be carefully considered by the practitioner, especially as many derived thresholds have no specific meaning regarding established AIS transceiving regulations or the maritime domain in general.
@@ 11 Conclusion
Maritime AIS data is inherently noisy due to technical inaccuracies, seafarer inattention, or lacking compliance. However, trajectories extracted from AIS messages are valuable for maritime surveillance, domain awareness, and algorithmic testing, for which we propose a data-driven, open-source trajectory extraction and refinement framework. The extraction process is based on a split-point procedure that uses quantile boundaries of five empirically derived distribution functions to decide whether two consecutive messages belong to the same trajectory or should be split. We found that this procedure effectively eliminates anomalies in the data without resorting to expert-derived thresholds, making it universally adaptable to any AIS data set. Nevertheless, carefully scrutinizing the quantile boundary and threshold values is recommended, as they may vary depending on data availability. For spatial understanding, we introduced the concept of average absolute change of course, $\bar{\Delta}(\cdot)$, which allows us to assess the presence or absence of erratic positional movement in a trajectory and recommended the number of messages per trajectory ($n_{\mathrm{msg}}$) as well as the convex hull area ($A^{C}$) as means to control $\bar{\Delta}(\cdot)$. The entire extraction and assessment pipeline, described by this article, is publicly available as an open-source Python package at Paulig (2024).